### v1

# Text-to-State Mapping for Non-Resolution Reasoning: A Formal Framework for Collapse Deferment in LLM Inference

Abstract
Large language models exhibit a systematic tendency toward early semantic commitment: given ambiguous input, they collapse multiple valid interpretations into a single response before sufficient context is available. This premature collapse discards information that may prove essential as dialogue evolves.
We present a formal framework for text-to-state mapping ($\phi:\mathcal{T}\to\mathcal{S}$) that transforms natural language into a non-collapsing state space where multiple interpretations coexist. The mapping decomposes into three stages: conflict detection, interpretation extraction, and state construction. We instantiate $\phi$ with a hybrid extraction pipeline that combines rule-based segmentation for explicit conflict markers (adversative conjunctions, hedging expressions) with LLM-based enumeration of implicit ambiguity (epistemic, lexical, structural).
On a test set of 68 ambiguous sentences, the resulting states preserve interpretive multiplicity: using hybrid extraction we obtain mean state entropy $H=1.087$ bits across ambiguity categories, compared to $H=0$ for collapse-based baselines that commit to a single interpretation. We additionally instantiate the rule-based conflict detector for Japanese markers (kedo, kamoshirenai, etc.) to illustrate cross-lingual portability of the conflict detection stage. This framework extends Non-Resolution Reasoning (NRR) by providing the missing algorithmic bridge between text and the NRR state space, enabling architectural collapse deferment in LLM inference.
© 2026 Kei Saito. Licensed under CC BY 4.0.
https://creativecommons.org/licenses/by/4.0/
Keywords: text-to-state mapping, non-resolution reasoning, collapse deferment, ambiguity preservation, LLM inference architecture
1 Introduction
1.1 The Early Semantic Commit Problem
Large language models (LLMs) have achieved remarkable capabilities in natural language understanding and generation. However, these systems share a fundamental architectural characteristic: they commit to interpretations early in the inference process, often before sufficient context is available to disambiguate.
Consider the input: “I want to quit my job, but I also don’t want to quit.” A standard LLM must produce a response, which requires committing to an interpretation of this ambivalent statement. Typical responses either:
  • Select one pole of the ambivalence as primary (“It sounds like you’re leaning toward staying…”)
  • Reframe the ambivalence as a problem requiring resolution (“Let’s make a pros and cons list…”)
  • Request clarification (“Which feeling is stronger?”)
Each of these responses presupposes that the ambivalence should be resolved. The system cannot maintain both interpretations as valid, coexisting states.
This early semantic commit pattern stems from architectural choices embedded in standard LLM design:
Softmax Normalization.
Attention mechanisms use softmax to produce probability distributions that sum to unity, forcing interpretations to compete for probability mass rather than coexist independently.
Autoregressive Generation.
Each token generation step commits irrevocably to a single choice. Uncertainty must be resolved before output, not deferred.
Training Objectives.
Cross-entropy loss rewards convergence toward ground-truth tokens, treating maintained ambiguity as prediction error rather than valid epistemic state.
1.2 Non-Resolution Reasoning
Non-Resolution Reasoning (NRR) [1] offers an alternative paradigm that treats ambiguity preservation as a valid computational mode. NRR introduced the state space $\mathcal{S}=\{(v_{i},c_{i},w_{i})\}$ where multiple interpretations coexist with independent activation weights, and demonstrated that prompt-based “NRR-lite” implementations maintain high entropy ($H=0.63$) at ambiguous turns while baseline models collapse ($H=0.10$).
However, a critical component remained undefined: how does text become state? The mapping $\phi:\mathcal{T}\to\mathcal{S}$ that transforms natural language input into the NRR state space was described conceptually but not algorithmically specified.
1.3 Contributions
This paper provides the missing algorithmic bridge. Our contributions are:
  1. 1.
    Formal Definition of $\phi$: We define the text-to-state mapping as a composition of three stages: conflict detection, interpretation extraction, and state construction.
  2. 2.
    Hybrid Extraction Algorithm: We combine rule-based segmentation (for explicit conflict markers) with LLM-based enumeration (for implicit ambiguity), providing both reproducibility and coverage.
  3. 3.
    Experimental Validation: We evaluate on 68 ambiguous sentences using three LLMs (GPT-4, Gemini, Claude), demonstrating consistent entropy preservation ($H=1.087$) across ambiguity categories.
  4. 4.
    Cross-Lingual Portability: We instantiate the rule-based conflict detector for Japanese markers, demonstrating that the framework is not tied to a single language.
1.4 Paper Organization
Section 2 reviews related work on uncertainty in LLM inference. Section 3 presents the formal framework for text-to-state mapping. Section 4 provides algorithms and implementation details. Section 5 reports experimental results. Section 6 briefly illustrates application to psychological support contexts. Section 7 discusses limitations and future work. Section 8 concludes.
2 Related Work
2.1 Uncertainty in LLM Inference
Uncertainty quantification in neural networks has received substantial attention [3, 4]. For LLMs specifically, calibration studies examine whether model confidence aligns with accuracy [5]. However, these approaches focus on measuring uncertainty rather than preserving it architecturally.
Epistemic uncertainty (model uncertainty) and aleatoric uncertainty (data uncertainty) are typically distinguished [6]. Our framework addresses a different dimension: interpretive multiplicity, where multiple valid readings coexist independent of model confidence.
2.2 Multi-Hypothesis Approaches
Beam search [7] and diverse decoding [8] maintain multiple candidates during generation but ultimately select a single output. Mixture-of-Experts architectures [9, 10] route inputs to specialized subnetworks but still produce collapsed outputs.
These approaches treat multiplicity as a means to better single outputs, not as a representational end state. NRR inverts this relationship: non-collapse is the goal, with singular output produced through non-destructive projection when required.
2.3 Ambiguity in NLP
Word sense disambiguation [11] and coreference resolution [12] address specific ambiguity types but assume disambiguation is always desirable. Pragmatic inference models [13] reason about speaker intent but still converge to single interpretations.
2.4 Relation to NRR Framework
This paper builds directly on the NRR framework [1]. Paper 1 established the state space formalism and demonstrated entropy preservation through prompt engineering. NRR defines eight operators ($\sigma,\alpha,\rho,\iota,\delta,\tau,\kappa,\pi$) for manipulating non-collapsing states (see Appendix D for summary).
This paper fills the gap between text and state space: without a defined $\phi$, the NRR operators have no input to operate on. We provide that input transformation.
3 Formal Framework
3.1 Preliminary Definitions
Definition 1 (Text Space).
Let $\mathcal{T}$ denote the space of natural language texts, where each $T\in\mathcal{T}$ is a sequence of tokens $(t_{1},t_{2},\ldots,t_{n})$.
Definition 2 (Semantic State Space).
Following [1], a semantic state space is a structure $\mathcal{S}=(V,C,W,M,\preceq)$ where:
  • $V\subseteq\mathbb{R}^{d}$ is a set of semantic vector representations
  • $C$ is a set of context identifiers
  • $W:V\times C\to[0,1]$ is an activation weight function
  • $M$ is a set of metadata flags (including conflict markers)
  • $\preceq$ is a partial order on $V$ (subsumption relation)
A state instance is a set $S=\{(v_{i},c_{i},w_{i},m_{i})\}_{i=1}^{n}$ drawn from this space. Intuitively, a state instance corresponds to a superposition of multiple interpretations, each tagged by its context, weight, and metadata—analogous to maintaining multiple hypotheses in parallel rather than committing to one.
Definition 3 (Interpretation).
An interpretation of text $T$ is a tuple $\iota=(v,c,w,m)$ where:
  • $v\in V$ is the semantic vector encoding one possible meaning
  • $c\in C$ identifies the context under which this interpretation holds
  • $w\in[0,1]$ is the initial activation weight
  • $m\in M$ contains metadata (conflict flags, source markers, etc.)
3.2 The Text-to-State Mapping
Definition 4 (Text-to-State Mapping).
The mapping $\phi:\mathcal{T}\to\mathcal{S}$ transforms input text $T$ to a state space instance:
$\phi(T)=\{(v_{i},c_{i},w_{i},m_{i})\mid i\in\mathcal{I}(T)\}$ (1)
where $\mathcal{I}(T)$ is the interpretation index set extracted from $T$.
The mapping $\phi$ decomposes into three sequential operations:
$\phi=\psi_{\text{state}}\circ\psi_{\text{interp}}\circ\psi_{\text{conflict}}$ (2)
where:
  1. 1.
    $\psi_{\text{conflict}}:\mathcal{T}\to\mathcal{T}\times\{0,1\}^{k}$ detects conflict markers
  2. 2.
    $\psi_{\text{interp}}:\mathcal{T}\times\{0,1\}^{k}\to\mathcal{I}$ extracts interpretations
  3. 3.
    $\psi_{\text{state}}:\mathcal{I}\to\mathcal{S}$ constructs the state space
3.3 Stage 1: Conflict Detection
The conflict detection stage identifies linguistic markers indicating multiple coexisting interpretations.
Definition 5 (Conflict Markers).
A conflict marker set $\mathcal{M}_{\text{conf}}$ is a collection of linguistic patterns indicating interpretive multiplicity:
$\mathcal{M}_{\text{conf}}=\mathcal{M}_{\text{explicit}}\cup\mathcal{M}_{\text{implicit}}\cup\mathcal{M}_{\text{structural}}$ (3)
where:
  • $\mathcal{M}_{\text{explicit}}$: Explicit contradiction markers (e.g., “but”, “however”, “on the other hand”, “yet”)
  • $\mathcal{M}_{\text{implicit}}$: Hedging and uncertainty markers (e.g., “maybe”, “perhaps”, “might”, “I think”)
  • $\mathcal{M}_{\text{structural}}$: Structural ambiguity indicators (e.g., “both X and Y”, “either… or”, coordination ambiguity)
The marker set extends to multilingual contexts; for Japanese, corresponding markers include adversatives (kedo, demo, shikashi), hedging expressions (kamoshirenai), and epistemic markers (to omou). See Appendix for the complete taxonomy.
Definition 6 (Conflict Detection Function).
The conflict detection function $\psi_{\text{conflict}}$ returns:
$\psi_{\text{conflict}}(T)=(T,\mathbf{f})$ (4)
where $\mathbf{f}=(f_{1},f_{2},\ldots,f_{k})\in\{0,1\}^{k}$ is a binary feature vector indicating presence of each marker type.
3.4 Stage 2: Interpretation Extraction
Interpretation extraction transforms text into a set of distinct meaning representations using a hybrid approach.
Definition 7 (Hybrid Extraction).
The interpretation extraction function combines two mechanisms:
$\psi_{\text{interp}}(T,\mathbf{f})=\text{Merge}(\psi_{\text{rule}}(T,\mathbf{f}),\psi_{\text{LLM}}(T,\mathbf{f}))$ (5)
Rule-Based Extraction.
For texts with detected conflict markers ($\|\mathbf{f}\|_{1}>0$), rule-based extraction segments text at conflict boundaries:
$\psi_{\text{rule}}(T,\mathbf{f})=\{(s_{j},c_{j})\mid j\in J(T,\mathbf{f})\}$ (6)
where $J(T,\mathbf{f})$ indexes segments delimited by conflict markers, $s_{j}$ is the $j$-th segment, and $c_{j}$ is the context label derived from marker type (e.g., “pre-adversative”, “post-adversative”, “hedge-scope”).
LLM-Based Extraction.
For complex or implicit ambiguity, LLM-based extraction prompts a language model to enumerate interpretations:
$\psi_{\text{LLM}}(T,\mathbf{f})=\text{LLM}_{\theta}(\text{Prompt}_{\text{interp}}(T,\mathbf{f}))$ (7)
where $\text{Prompt}_{\text{interp}}$ is a structured prompt requesting exhaustive interpretation enumeration.
Merge Operation.
The merge function removes duplicates via semantic similarity:
$\text{Merge}(I_{1},I_{2})=I_{1}\cup\{i\in I_{2}\mid\forall j\in I_{1}:\text{sim}(i,j)<\tau\}$ (8)
where $\text{sim}(i,j)$ is cosine similarity computed over a sentence-level embedding space (e.g., Sentence-BERT [2]), and $\tau\in(0,1)$ is a deduplication threshold (typically $\tau=0.85$).
3.5 Stage 3: State Construction
Definition 8 (State Construction).
Given interpretations $\mathcal{I}=\{(\text{interp}_{i},\text{ctx}_{i},\text{conf}_{i})\}$, the state construction function produces:
$\psi_{\text{state}}(\mathcal{I})=\{(v_{i},c_{i},w_{i},m_{i})\}_{i=1}^{|\mathcal{I}|}$ (9)
where:
$v_{i}$ $=\text{Embed}(\text{interp}_{i})\in\mathbb{R}^{d}$ (10)
$c_{i}$ $=\text{ContextEncode}(\text{ctx}_{i})\in C$ (11)
$w_{i}$ $=\text{conf}_{i}\cdot(1+\beta\cdot\mathbf{1}[\text{conflict}])$ (12)
$m_{i}$ $=\{\text{source}:\text{rule/LLM},\text{conflict}:\mathbf{f}\}$ (13)
The weight adjustment factor $\beta>0$ increases activation for interpretations from conflicting contexts, reflecting the NRR principle that conflict indicates information density.
3.6 Information-Theoretic Properties
Definition 9 (State Entropy).
For a state $S=\{(v_{i},c_{i},w_{i},m_{i})\}_{i=1}^{n}$, the state entropy is the Shannon entropy of the normalized weight distribution:
[…]
Normalization is applied solely for entropy computation; the weights $w_{i}$ themselves remain unnormalized in the state representation.
Theorem 1 (Non-Collapse of $\phi$).
For any text $T$ with $|\mathcal{I}(T)|>1$ distinct extractable interpretations, the mapping $\phi$ produces a state $S$ such that:
  1. 1.
    (Structural non-collapse) $|S|\geq 2$, i.e., at least two interpretations are preserved in the state space.
  2. 2.
    (Entropy preservation) The state entropy satisfies $H(S)>H_{\text{baseline}}(T)$ where $H_{\text{baseline}}(T)\approx 0$ is the entropy of a collapse-based model that commits to a single interpretation.
Proof.
(1) By construction, $\psi_{\text{interp}}$ extracts at least two interpretations when conflict markers are detected or when LLM enumeration returns multiple candidates. The merge operation removes only near-duplicates, preserving distinct interpretations.
(2) For collapse-based models, $p^{*}=1$ for the selected interpretation, yielding $H_{\text{baseline}}=0$. For $\phi$, with $|S|\geq 2$ and all $w_{i}>0$, the distribution has support on multiple elements, guaranteeing $H(S)>0$. ∎
Remark 1.
The theorem establishes a relative guarantee: $\phi$ preserves more information than collapse-based alternatives. The absolute entropy depends on extraction quality and weight calibration, addressed empirically in Section 5.
4 Algorithm and Implementation
4.1 Text-to-State Mapping Algorithm
Algorithm 1 Text-to-State Mapping ($\phi$)
0: Text $T$, embedding function Embed, LLM $\mathcal{L}_{\theta}$
0: State $S=\{(v_{i},c_{i},w_{i},m_{i})\}$
1:// Stage 1: Conflict Detection
[…]
4:// Stage 2: Interpretation Extraction
[…]
6:if has_conflict then
7:  $I_{\text{rule}}\leftarrow\text{SegmentAtMarkers}(T,\mathbf{f})$
8:end if
[…]
11:// Stage 3: State Construction
[…]
13:for $(\text{interp},\text{ctx},\text{conf})\in I$ do
14:  $v\leftarrow\text{Embed}(\text{interp})$
15:  $c\leftarrow\text{ContextEncode}(\text{ctx})$
16:  $w\leftarrow\text{conf}\times(1+\beta\cdot\mathbf{1}[\text{has\_conflict}])$
17:  $m\leftarrow\{\text{conflict}:\mathbf{f},\text{source}:\text{GetSource}(\text{interp})\}$
18:  $S\leftarrow S\cup\{(v,c,w,m)\}$
19:end for
20:return $S$
4.2 NRR Processing Pipeline
The mapping $\phi$ feeds into the NRR operator pipeline (Appendix D):
Algorithm 2 NRR Processing Pipeline
0: Text $T$, previous state $S_{\text{prev}}$ (or $\emptyset$)
0: Processed state $S^{\prime}$, output $o$
[…]
8:return $(S^{\prime},o)$
4.3 Computational Complexity
  • Conflict detection: $O(|T|\times|\mathcal{M}|)$ where $|\mathcal{M}|$ is the marker vocabulary size
  • Rule-based extraction: $O(|T|)$ for segmentation
  • LLM-based extraction: Dominated by LLM inference cost
  • State construction: $O(|\mathcal{I}|\times d)$ for embedding
  • Merge deduplication: $O(|\mathcal{I}|^{2})$ for pairwise similarity
For typical inputs ($|T|<100$ tokens, $|\mathcal{I}|<10$ interpretations), the overhead is negligible compared to LLM inference.
5 Experimental Validation
5.1 Research Questions
  • RQ1: Does $\phi$ preserve interpretive multiplicity for ambiguous inputs?
  • RQ2: Which ambiguity categories are captured by rule-based vs. LLM-based extraction?
  • RQ3: Does the framework generalize across languages (English, Japanese)?
  • RQ4: How consistent are results across different LLMs (GPT-4, Gemini, Claude)?
5.2 Test Set
We constructed a test set of 68 ambiguous sentences. For rule-based extraction (adversative, hedging), we include both English and Japanese sentences to demonstrate cross-lingual portability of $\psi_{\text{conflict}}$. For LLM-based extraction (epistemic, lexical, structural), we use English only to isolate the contribution of our framework from the multilingual capabilities of the underlying LLMs.
Category Total EN JP Method
Adversative 20 10 10 Rule-based
Hedging 20 10 10 Rule-based
Epistemic 8 8 LLM-based
Lexical 10 10 LLM-based
Structural 10 10 LLM-based
Total 68 48 20
Table 1: Test set composition. Japanese sentences are included only for rule-based categories to demonstrate cross-lingual portability of conflict detection.
Figure 1: Test set composition across five ambiguity categories (68 sentences total). Rule-based extraction targets adversative and hedging categories with both English and Japanese sentences. LLM-based extraction targets epistemic, lexical, and structural categories with English sentences only.
Adversative sentences contain explicit contradiction markers (“but”, “however”, Japanese kedo, demo). Example: “I want to quit my job, but I also don’t want to quit.”
Hedging sentences contain uncertainty markers (“maybe”, “perhaps”, Japanese kamoshirenai). Example: “Maybe I should apply for that position.”
Epistemic sentences contain belief markers (“I think”, “I believe”). Example: “I think I made the right choice.”
Lexical ambiguity sentences contain polysemous words. Example: “I saw her duck.”
Structural ambiguity sentences have multiple parse trees. Example: “I shot an elephant in my pajamas.”
5.3 Metrics
  • State Size $|S|$: Number of interpretations in the state space
  • State Entropy $H(S)$: Shannon entropy of the weight distribution
  • Conflict Detection Rate: Proportion of sentences where markers were detected
  • EPR (Entropy Preservation Ratio): $H(S)/H_{\max}$ where $H_{\max}=\log_{2}|S|$
5.4 Results: Rule-Based Extraction
We first evaluate rule-based extraction alone. Table 2 shows that rule-based methods effectively capture adversative and hedging patterns ($H\approx 1.0$) but fail on epistemic, lexical, and structural ambiguity ($H=0$).
Category $|S|$ $H(S)$ Method
Adversative 2.10 1.037 Rule-based
Hedging 2.05 1.002 Rule-based
Epistemic 1.00 0.000 Rule-based
Lexical 1.00 0.000 Rule-based
Structural 1.00 0.000 Rule-based
Table 2: Rule-based extraction results. Epistemic, lexical, and structural categories require LLM-based extraction.
5.5 Results: LLM-Based Extraction
To address the limitations of rule-based extraction, we evaluated LLM-based interpretation extraction using three models: GPT-4, Gemini, and Claude. Each model was prompted to enumerate all possible interpretations for 28 English sentences (8 epistemic + 10 lexical + 10 structural).
Category GPT-4 Gemini Claude Mean
Epistemic $|S|$ 3.00 3.00 3.75 3.25
Epistemic $H(S)$ 1.577 1.558 1.885 1.673
Lexical $|S|$ 2.00 2.30 2.20 2.17
Lexical $H(S)$ 0.990 1.024 1.008 1.007
Structural $|S|$ 2.20 2.10 2.10 2.13
Structural $H(S)$ 1.056 0.821 1.027 0.968
Table 3: LLM-based extraction results (English only) across three models. All categories achieve $H>0$, with epistemic showing highest entropy due to multiple interpretive layers (assertion vs. hedge vs. self-justification).
Figure 2: LLM-based extraction results across three models (GPT-4, Gemini, Claude) for English sentences. The dashed line indicates $H_{\max}=1.0$ for binary interpretations. All models achieve $H>0$ across all categories, with epistemic showing the highest entropy due to multiple interpretive layers.
Key Findings.
LLM-based extraction successfully recovers interpretive multiplicity for all categories that rule-based methods failed on:
  • Epistemic: Mean $|S|=3.25$, $H=1.673$ (highest entropy, reflecting multiple interpretive layers)
  • Lexical: Mean $|S|=2.17$, $H=1.007$ (captures polysemy)
  • Structural: Mean $|S|=2.13$, $H=0.968$ (captures parse ambiguity)
Claude extracted the most interpretations on average ($|S|=3.75$ for epistemic), while all three models achieved $H>0$ for 100% of test sentences.
5.6 Combined Results
Table 4 presents the final results combining rule-based extraction (for adversative and hedging, including Japanese sentences) with LLM-based extraction (for epistemic, lexical, and structural, English only).
Category N $|S|$ $H(S)$ Method
Adversative 20 2.10 1.037 Rule (EN+JP)
Hedging 20 2.05 1.002 Rule (EN+JP)
Epistemic 8 3.25 1.673 LLM (EN)
Lexical 10 2.17 1.007 LLM (EN)
Structural 10 2.13 0.968 LLM (EN)
Overall 68 2.24 1.087 Hybrid
Table 4: Combined results using hybrid extraction. All categories achieve $|S|\geq 2$ and $H>0$, confirming Theorem 1.
Figure 3: Combined experimental results. (A) State entropy $H(S)$ by category, showing rule-based extraction (teal) achieves $H\approx 1.0$ for adversative and hedging, while LLM-based extraction (coral) captures epistemic, lexical, and structural ambiguity. (B) Number of interpretations extracted per category. (C) Overall comparison: baseline systems collapse to $H=0$, while $\phi$ preserves $H=1.087$ across all 68 sentences.
5.7 Baseline Comparison
For comparison, a collapse-based system that commits to a single interpretation yields $H_{\text{baseline}}=0$ for all inputs. Note that this baseline entropy is a theoretical value derived from the architectural constraint that standard LLMs maintain a single active interpretation during inference; we do not measure $H$ empirically for baseline systems, as they do not expose internal interpretive states. The improvement from $\phi$ is:
$\Delta H=H(\phi(T))-H_{\text{baseline}}=1.087-0=+1.087\text{ bits}$ (15)
This represents substantial information preservation. For epistemic sentences specifically, $\Delta H=1.673$ bits—exceeding the theoretical maximum for binary states ($H_{\max}=1.0$) because LLMs extract 3+ interpretations on average.
Answering the Research Questions.
(RQ1) Yes, $\phi$ preserves interpretive multiplicity: mean $|S|=2.24$ and $H=1.087$ across all categories, compared to $|S|=1$ and $H=0$ for baselines.
(RQ2) Rule-based extraction captures explicit markers (adversative, hedging) with $H\approx 1.0$. LLM-based extraction is required for implicit ambiguity (epistemic, lexical, structural), achieving H=0.97H=0.971.671.67.
(RQ3) The rule-based conflict detector processes Japanese markers (kedo, demo, kamoshirenai) equivalently to English, demonstrating cross-lingual portability of $\psi_{\text{conflict}}$.
(RQ4) All three LLMs (GPT-4, Gemini, Claude) achieve $H>0$ for 100% of English test sentences. Claude extracts the most interpretations (mean $|S|=3.75$ for epistemic), while Gemini is more conservative ($|S|=3.00$). Despite variation in $|S|$, all models successfully preserve interpretive multiplicity.
5.8 Limitations
While the experimental results validate the framework, several limitations remain:
  • Test set size: 68 sentences across 5 categories provides initial validation but larger-scale evaluation is needed.
  • LLM variability: Different models extract different numbers of interpretations (Claude: 3.75, Gemini: 3.00 for epistemic), suggesting sensitivity to model choice.
  • Human evaluation: We measured entropy but did not evaluate whether extracted interpretations align with human judgments.
  • Weight calibration: Confidence weights from LLMs may not accurately reflect interpretation plausibility.
  • Cross-lingual scope: We demonstrate Japanese support only for rule-based extraction; full multilingual evaluation requires further work.
6 Illustrative Application: Psychological Support
We briefly illustrate how $\phi$ applies to psychological support contexts, where ambivalence is common and premature resolution may be counterproductive.
6.1 Motivation
In therapeutic contexts, statements like “I love them, but being with them hurts” express genuine ambivalence that clients benefit from exploring rather than resolving. Standard LLMs tend to:
  • Diagnose the ambivalence as a “problem”
  • Offer resolution strategies (pros/cons lists, decision frameworks)
  • Select one pole as primary
With $\phi$, the ambivalent statement maps to a state with two interpretations:
$S=\{$ $(\text{``I love them''},\text{pre-adv},0.96,\{\text{conflict}:\text{True}\}),$
$(\text{``being with them hurts''},\text{post-adv},0.96,\{\text{conflict}:\text{True}\})\}$
Both interpretations are preserved with equal weight, enabling responses that acknowledge the coexistence rather than forcing resolution.
6.2 Limitations
This application is illustrative, not validated. Whether NRR-based systems produce therapeutically beneficial outcomes requires:
  • Clinical trials with appropriate ethical oversight
  • Comparison with established therapeutic approaches
  • Long-term outcome measurement
NRR is a computational architecture, not a therapeutic intervention.
7 Discussion
7.1 Implications for LLM Architecture
Natural language is inherently underspecified: identical lexical forms do not guarantee identical referential states due to perspectival variation, temporal–spatial indexicality, and context-dependent conceptualization. Consequently, committing to a single semantic interpretation at inference time is not an information-gain event but an information-loss event. In this view, the role of $\phi$ is not merely to enumerate alternative readings, but to formalize a separation between the factual layer (what was expressed) and the interpretive layer (what it could plausibly denote), allowing systems to defer semantic collapse until contextual constraints warrant resolution.
The text-to-state mapping demonstrates that early semantic commitment is a design choice, not an architectural necessity. Systems can:
  1. 1.
    Detect interpretive multiplicity at input
  2. 2.
    Maintain multiple interpretations through processing
  3. 3.
    Defer collapse until external output is required
  4. 4.
    Preserve internal state after output (non-destructive projection)
This has implications for dialogue systems, where context evolves across turns and early commitments may prove incorrect.
7.2 Distinction from Prompt-Based Enumeration
A natural question arises: if LLMs can enumerate multiple interpretations when prompted, why is $\phi$ necessary? Note that prompting an LLM to enumerate interpretations results only in a textual list, not in a persistent or manipulable state representation. The model generates these interpretations sequentially and does not maintain them as simultaneously active hypotheses. In contrast, $\phi$ constructs an explicit state $S$ that can be preserved and transformed by NRR operators across dialogue turns. The distinction is between generating ambiguity (textual enumeration) and maintaining ambiguity (state persistence).
7.3 Limitations
Marker Coverage.
The conflict marker taxonomy is not exhaustive. Languages beyond English and Japanese require additional marker sets.
LLM Dependence.
LLM-based extraction depends on external model availability and introduces non-determinism. Rule-based extraction is reproducible but limited in coverage.
Weight Calibration.
Current weight assignment ($w=\text{conf}\times(1+\beta)$) is heuristic. Learned weight calibration could improve entropy properties.
Evaluation Scope.
The 68-sentence test set covers common ambiguity patterns but may not represent all real-world inputs.
7.4 Future Work
  • Native Non-Collapsing Inference: Integrate $\phi$ into LLM architectures rather than as a preprocessing step
  • Learned Conflict Detection: Train classifiers for marker detection rather than pattern matching
  • Cross-Lingual Extension: Develop marker taxonomies for additional languages
  • Human Evaluation: Assess perceived quality of non-collapsing responses in user studies
8 Conclusion
We have presented a formal framework for text-to-state mapping ($\phi:\mathcal{T}\to\mathcal{S}$) that transforms natural language into a non-collapsing state space. The mapping decomposes into conflict detection, interpretation extraction, and state construction, with a hybrid approach combining rule-based and LLM-based extraction.
Experimental evaluation on 68 ambiguous sentences demonstrates that the mapping preserves interpretive multiplicity: mean state entropy $H=1.087$ bits across all categories, compared to $H=0$ for collapse-based baselines.
The framework provides the missing algorithmic component of Non-Resolution Reasoning: the transformation from text to the state space on which NRR operators act. This enables architectural collapse deferment in LLM inference, supporting applications where ambiguity preservation is valuable.
Early semantic commitment is a design choice. This paper shows how to choose otherwise.
Appendix A Conflict Marker Taxonomy
Category English Japanese (romanized) Context Label
Explicit Contradiction
Adversative but, however, yet, although kedo, demo, shikashi, daga pre-adv, post-adv
Contrastive on the other hand, whereas ippou de, hanmen contrast-A, contrast-B
Concessive even though, despite nimo kakawarazu concede, main
Implicit Uncertainty
Hedging maybe, perhaps, might kamoshirenai, tabun hedge-scope
Epistemic I think, I believe, it seems to omou, ki ga suru epistemic-stance
Modal could, would, should beki, hazu, darou modal-world
Structural Ambiguity
Coordination both X and Y, either X or Y mo…mo, ka…ka coord-A, coord-B
Scope all…not, every…some wide-scope, narrow-scope
Table 5: Conflict marker taxonomy (English and Japanese)
Appendix B LLM Prompt Template
The following prompt template is used for LLM-based interpretation extraction:
Given the text: "[T]"
[If conflict markers detected:
"Note: This text contains potential ambiguity markers."]
List ALL possible interpretations as distinct meanings.
For each interpretation, provide:
1. The interpretation (a clear restatement of one possible meaning)
2. The context/condition under which this interpretation holds
3. Confidence weight from 0.0 to 1.0
Format each as:
INTERP: [interpretation]
CONTEXT: [context]
CONFIDENCE: [0.0-1.0]
Appendix C Sample Outputs
This appendix provides representative examples of how $\phi$ processes different ambiguity types. We show one example each from lexical ambiguity (LLM-based), structural ambiguity (LLM-based), and adversative (rule-based with Japanese).
Example 1: Lexical Ambiguity (LLM-based)
Input: “I saw her duck.”
GPT-4 Output:
  "id": "lex_en_01",
  "interpretations": [
      "meaning": "I observed her pet duck (the bird).",
      "context": "noun-reading",
      "confidence": 0.5
      "meaning": "I saw her lower her head quickly.",
      "context": "verb-reading",
      "confidence": 0.5
State constructed by $\phi$: $|S|=2$, $H(S)=1.00$ bits
Example 2: Structural Ambiguity (LLM-based)
Input: “I shot an elephant in my pajamas.”
GPT-4 Output:
  "id": "str_en_01",
  "interpretations": [
      "meaning": "The speaker was wearing pajamas
                  when they shot the elephant.",
      "context": "PP attaches to subject",
      "confidence": 0.7
      "meaning": "The elephant was inside the speaker’s
                  pajamas when it was shot.",
      "context": "PP attaches to object",
      "confidence": 0.3
State constructed by $\phi$: $|S|=2$, $H(S)=0.88$ bits
Example 3: Japanese Adversative (Rule-based)
Input: “Yametai kedo yametakunai” (“I want to quit, but I also don’t want to quit.”)
Rule-based segmentation: Marker “kedo” detected.
Segments extracted:
  • Segment 1: “yametai” (“I want to quit”) — context: pre-adv
  • Segment 2: “yametakunai” (“I don’t want to quit”) — context: post-adv
Weights: Both segments receive $w=0.5$ (equal weight for adversative).
State constructed by $\phi$: $|S|=2$, $H(S)=1.00$ bits
These examples illustrate the difference between LLM-based extraction (which enumerates semantic interpretations) and rule-based extraction (which segments based on syntactic markers). Both approaches yield states with $H>0$, preserving interpretive multiplicity.
Appendix D NRR Operator Summary
NRR uses eight operators to transform the state space $\mathcal{S}$, as introduced in [1]. We summarize their functions here for reference.
  • $\sigma$ (Stripping): Removes learned biases inducing collapse.
  • $\alpha$ (Label-Free Abstraction): Projects to geometric relations without labels.
  • $\rho$ (Positioning): Assigns contextual coordinates (implements CIT).
  • $\iota$ (Invariance): Identifies structure stable across contexts.
  • $\delta$ (Dampening): Reduces activation differentials.
  • $\tau$ (Deferred Resolution): Postpones projection to output boundary.
  • $\kappa$ (CPP Integration): Preserves contradictory states with conflict tags.
  • $\pi$ (Persistence): Maintains state across turns with decay.
The composition $\pi\circ\kappa\circ\tau\circ\delta\circ\iota\circ\rho\circ\alpha\circ\sigma$ defines the full pipeline, with $\phi$ (this paper) providing the input transformation from text to initial state.
Acknowledgments
The author gratefully acknowledges the assistance of large language models—including Claude (Anthropic), ChatGPT (OpenAI), and Gemini (Google)—for linguistic refinement, structural feedback, and LaTeX formatting during the preparation of this manuscript. All theoretical concepts, formal definitions, and conclusions are solely the responsibility of the author.